# hep-ph/***

### v1

@@ 4.4 Results for Case 2)
In figs. 9 and 10 we provide results of the numerical scans for $t$ even corresponding to $u$ even ($s=1$, $t=2$ and thus $u=0$) and for $t$ odd/$u$ odd ($s=1$, $t=1$, meaning $u=1$), respectively. The colour-coding is the same as in fig. 3 and the definition of the quantity $\overline{y}$ can be found in eq. (75). We see that the same orders of magnitude for the BRs and CR are predicted as in Case 1) and that also for Case 2) the prospective limits for both processes $\mu\to 3\,e$ and $\mu-e$ conversion in aluminium can be reached, while $\mathrm{BR}(\mu\to e\gamma)$ is expected to be smaller than $6\times 10^{-16}$ in the portion of parameter space that is allowed. Furthermore, this part of the parameter space is characterised by $\mu_{0}$ being larger than $2\,\mbox{keV}$, whereas the scanned intervals of $M_{0}$ and $\theta_{R}$ are not constrained. Nevertheless, we also observe certain differences compared to the results for Case 1). In particular, both BRs reveal lower limits, $\mathrm{BR}(\mu\to e\gamma)\gtrsim 10^{-19}$ and $\mathrm{BR}(\mu\to 3\,e)\gtrsim 10^{-21}$, and the CR is predicted to be larger than $5\times 10^{-22}$ for values of $M_{0}$ which are not too close to $M_{0}\approx 6.5\,\mbox{TeV}$, where the cancellation occurs.161616 We actually exclude all values of the CR which are obtained for $4.5\,\mbox{TeV}\lesssim M_{0}\lesssim 7.5\,\mbox{TeV}$. As expected, for $t$ even, see fig. 9, no dependence of the flavour observables on the angle $\theta_{R}$ is seen. Consequently, we observe less spread in the data points. For $t$ odd, see fig. 10, a clear dependence on $\theta_{R}$ is observed. In contrast to the results for Case 1), the BRs and CR are now enhanced for $\sin 2\,\theta_{R}$ close to zero, since then at least one Yukawa coupling becomes large (compare the results found in [47]), while for $|\sin 2\,\theta_{R}|$ being maximal the Yukawa couplings and, hence, the size of the cLFV signals are suppressed. This behaviour is very similar for light neutrino masses with IO and $m_{0}=0.015\,\mbox{eV}$ as well as in situations in which the lightest neutrino mass $m_{0}$ is set to zero.
@@ 4.5 Results for Case 3 a)
giving rise to $\sin^{2}\theta_{13}\approx 0.022$ and $\sin^{2}\theta_{23}\approx 0.61$. The solar mixing angle is fixed by the choice of the CP symmetry, i.e. the parameter $s$, and the free angle and all values of $s$, $0\leq s\leq n-1$, allow for a good fit of the lepton mixing angles, usually for two different values of the free angle. For examples of viable fits, see [33] and more recently [47].
@@
For the choice $m$ even and $s$ odd (for $n=34$) we expect the results to depend on $\theta_{R}$ and this dependence should be similar to the one found for Case 1), compare [47]. As can be observed in fig. 13, the dependence is much milder (for light neutrino masses with NO and $m_{0}=0.03\,\mbox{eV}$) than in the other studied cases, Case 1) (see fig. 3), Case 2) for $t$ odd (see fig. 10) and Case 3 b.1) for $m$ even and $s$ odd or vice versa (see fig. 19). We have confirmed that for $m$ odd and $s$ even (employing $n=16$) a very similar result is obtained. Assuming light neutrino masses with IO and $m_{0}=0.015\,\mbox{eV}$ also leads to plots similar to those shown in fig. 13. For NO light neutrino masses and $m_{0}$ being very small ($m_{0}=0$ or $m_{0}=10^{-5}\,\mbox{eV}$) we find instead a larger dependence on $\theta_{R}$ like for Case 1) in fig. 3, when considering e.g. $n=16$, $m=1$ and $s=2$. Regarding the experimental prospects to detect cLFV signals involving the muon and the electron for Case 3 a) and $m$ even and $s$ odd or vice versa, these are similar to those for $m$ and $s$ both even or both odd, namely for the studied examples we obtain $4\times 10^{-21}\lesssim\mathrm{BR}(\mu\to e\gamma)\lesssim 6\times 10^{-16}$, $\mathrm{BR}(\mu\to 3\,e)\gtrsim 10^{-22}$ and $\mathrm{CR}(\mu-e,\mathrm{Al})\gtrsim 10^{-21}$, if the cancellation region, see footnote 16, is avoided.
Lastly, we would like to compare the different choices of $n$, $m$ and $s$ that we have analysed numerically. Like for Case 2), see fig. 11, we make use of averages corresponding to a certain binning of the data in the parameter $\mu_{0}$. We find that the results of the BRs and CR are nearly independent of the choice of $n$, $m$ and $s$, as can be clearly seen in fig. 14, where we show the results for light neutrino masses with NO and non-zero $m_{0}$. This behaviour can be explained with the fact that no enhancement of the BRs and CR is observed for certain values of the angle $\theta_{R}$, compare fig. 13. The corresponding plots for light neutrino masses with IO and $m_{0}=0.015\,\mbox{eV}$ reveal the same features, as expected from the above discussion.
Figure 15: Case 3 b.1), $m$ and $s$ both even. Predictions for $\mathrm{BR}(\mu\to e\gamma)$, $\mathrm{BR}(\mu\to 3\,e)$ and $\mathrm{CR}(\mu-e,\mathrm{Al})$ in the $\frac{s}{n}-\theta_{L}$-plane in the upper, middle and lower row. The group theory parameters are chosen as $n=20$ and $m=10$. The conventions are like in fig. 5 of Case 2) and $t$ even ($u$ even). Plots reflecting the features of examples with $m$ and $s$ both odd can be found in figs. 16 and 17, since these are displayed for the choice $\theta_{R}=0$.
@@
Figure 16: Case 3 b.1), $m$ odd, $s$ even and $\theta_{R}=0$. Predictions for $\mathrm{BR}(\mu\to e\gamma)$, $\mathrm{BR}(\mu\to 3\,e)$ and $\mathrm{CR}(\mu-e,\mathrm{Al})$ in the $\frac{s}{n}-\theta_{L}$-plane in the upper, middle and lower row. The group theory parameters are chosen as $n=20$ and $m=9$. For the remaining conventions see fig. 15. The dependence on $\theta_{R}$ is illustrated with six different values of $\theta_{R}$ in different colours in fig. 22 in appendix D. Results for an example of the combination $m$ even and $s$ odd and different values of $\theta_{R}$ are found in fig. 24 in the same appendix.
@@
We begin the analysis of the BRs and CR by evaluating their sizes in the $\frac{s}{n}-\theta_{L}$-plane [^19] The ratio $\frac{s}{n}$ is taken to be continuous, $0\leq\frac{s}{n}\leq 1$, and $s$ is either assumed to be even or odd. for a fixed ratio of $\frac{m}{n}$, $\frac{m}{n}=\frac{9}{20}$, $\frac{m}{n}=\frac{10}{20}$ and $\frac{m}{n}=\frac{11}{20}$, as well as for fixed values of the two scales $\mu_{0}$ and $M_{0}$, $\mu_{0}=1\,\mbox{keV}$ and $M_{0}=3\,\mbox{TeV}$, see figs. 15-17. The colour-coding in these figures is the same as in fig. 5 for Case 2). In particular, the dark (light) grey regions in the plots for $\mathrm{CR}(\mu-e,\mathrm{Al})$ indicate the parameter space in which the lepton mixing angles are accommodated to a certain degree, i.e. $\chi^{2}\leq 100\,(300)$. These regions are found to be consistent with those shown in corresponding plots in [33, 47]. The plots in the left in figs. 15-17 assume that light neutrino masses follow NO with $m_{0}=0.03\,\mbox{eV}$, while the plots in the right in these figures are based on an IO light neutrino mass spectrum with $m_{0}=0.015\,\mbox{eV}$. Fig. 15 represents the combination $m$ and $s$ both even and, thus, no dependence on $\theta_{R}$ is expected, and also observed, while in figs. 16 and 17 $m$ is odd and $s$ even such that the results depend on $\theta_{R}$. Here, they are shown for $\theta_{R}=0$, for which the matrix combination in square brackets in eq. (26) is diagonal. We complete the set of plots by figs. 22-24 in appendix D in which we display for $\frac{m}{n}=\frac{9}{20}$ and $\frac{m}{n}=\frac{11}{20}$ and $s$ even as well as for $\frac{m}{n}=\frac{10}{20}$ and $s$ odd the results for six different values of the angle $\theta_{R}$ in different colours, similar to e.g. fig. 6 for Case 2) and $t$ odd. We note that the results for $\frac{m}{n}=\frac{9}{20}$ ($\frac{m}{n}=\frac{11}{20}$) and $s$ even agree for the choice $\theta_{R}=0$ with those for $\frac{m}{n}=\frac{9}{20}$ ($\frac{m}{n}=\frac{11}{20}$) and $s$ odd and, similarly, the choice $\frac{m}{n}=\frac{10}{20}$ and $s$ odd with $\theta_{R}=0$ reproduces the results of $\frac{m}{n}=\frac{10}{20}$ and $s$ even. Consequently, the results for all possible combinations of $m$ and $s$ even and odd are covered with the six figures, figs. 15-17 and figs. 22-24. We see for $m$ and $s$ both even or both odd as well as for $m$ even and $s$ odd or vice versa and $\theta_{R}=0$ that in most situations the prospective bound on $\mathrm{BR}(\mu\to e\gamma)$ is passed in the entire parameter space – at least at the $3\,\sigma$ level, whereas the limit expected from Mu3E Phase 1 excludes about half of the parameter space at the $3\,\sigma$ level and the future bound from Mu3E Phase 2 considerably reduces the viable regions of parameter space. Employing the expected limits from COMET and Mu2e only leaves a small portion of parameter space in the $\frac{s}{n}-\theta_{L}$-plane allowed, which in general is hardly compatible with the regions preferred by the measured values of the lepton mixing angles. In addition, we observe that the future bounds are slightly more likely to be satisfied for light neutrino masses with NO and $m_{0}=0.03\,\mbox{eV}$ than in the case of an IO light neutrino mass spectrum with $m_{0}=0.015\,\mbox{eV}$ (compare plots in the left with those in the right in figs. 15-17). When considering the dependence on $\theta_{R}$ for the combinations $m$ even and $s$ odd or vice versa, see especially figs. 22-24 in appendix D, we note that the shape and size of the parameter space passing a certain future bound depends on the actual value of $\theta_{R}$ and for the larger ones among the shown values it is usually smaller. Furthermore, the regions of parameter space that lead to an agreement of the lepton mixing angles with experimental data shift as function of $\theta_{R}$. It is, thus, necessary to investigate each value of $\theta_{R}$ separately in order to make more quantitative statements about the constraints imposed by the future limits on the different $\mu-e$ transitions and their compatibility with the restrictions coming from fitting the lepton mixing angles well.
Figure 17: Case 3 b.1), $m$ odd, $s$ even and $\theta_{R}=0$. Predictions for $\mathrm{BR}(\mu\to e\gamma)$, $\mathrm{BR}(\mu\to 3\,e)$ and $\mathrm{CR}(\mu-e,\mathrm{Al})$ in the $\frac{s}{n}-\theta_{L}$-plane in the upper, middle and lower row. The group theory parameters are chosen as $n=20$ and $m=11$. For the remaining conventions see fig. 15. The dependence on $\theta_{R}$ is illustrated with six different values of $\theta_{R}$ in different colours in fig. 23 in appendix D. Results for an example of the combination $m$ even and $s$ odd and different values of $\theta_{R}$ are found in fig. 24 in the same appendix.
@@
Again, we have fixed the two scales $\mu_{0}$ and $M_{0}$ to $\mu_{0}=1\,\mbox{keV}$ and $M_{0}=3\,\mbox{TeV}$, respectively. From these estimates we see that the future bound on $\mathrm{BR}(\mu\to e\gamma)$ is usually passed as well as the limit on $\mathrm{BR}(\mu\to 3\,e)$ expected from Mu3E Phase 1, whereas Mu3E Phase 2 can exclude this scenario and also the experiments COMET and Mu2e have this potential, since $\mathrm{CR}(\mu-e,\mathrm{Al})$ is larger than $2\times 10^{-16}$ for the chosen values of the scales $\mu_{0}$ and $M_{0}$ and the assumed light neutrino mass spectra.

### v2

@@ 4.4 Results for Case 2)
In figs. 9 and 10 we provide results of the numerical scans for $t$ even corresponding to $u$ even ($s=1$, $t=2$ and thus $u=0$) and for $t$ odd/$u$ odd ($s=1$, $t=1$, meaning $u=1$), respectively. The colour-coding is the same as in fig. 3 and the definition of the quantity $\overline{y}$ can be found in eq. (75). We see that the same orders of magnitude for the BRs and CR are predicted as in Case 1) and that also for Case 2) the prospective limits for both processes $\mu\to 3\,e$ and $\mu-e$ conversion in aluminium can be reached, while $\mathrm{BR}(\mu\to e\gamma)$ is expected to be smaller than $6\times 10^{-16}$ in the portion of parameter space that is allowed. Furthermore, this part of the parameter space is characterised by $\mu_{0}$ being larger than $2\,\mbox{keV}$, whereas the scanned intervals of $M_{0}$ and $\theta_{R}$ are not constrained. Nevertheless, we also observe certain differences compared to the results for Case 1). In particular, both BRs reveal lower limits, $\mathrm{BR}(\mu\to e\gamma)\gtrsim 10^{-19}$ and $\mathrm{BR}(\mu\to 3\,e)\gtrsim 10^{-21}$, and the CR is predicted to be larger than $5\times 10^{-22}$ for values of $M_{0}$ which are not too close to $M_{0}\approx 6.5\,\mbox{TeV}$, where the cancellation occurs.202020 We actually exclude all values of the CR which are obtained for $4.5\,\mbox{TeV}\lesssim M_{0}\lesssim 7.5\,\mbox{TeV}$. As expected, for $t$ even, see fig. 9, no dependence of the flavour observables on the angle $\theta_{R}$ is seen. Consequently, we observe less spread in the data points. For $t$ odd, see fig. 10, a clear dependence on $\theta_{R}$ is observed. In contrast to the results for Case 1), the BRs and CR are now enhanced for $\sin 2\,\theta_{R}$ close to zero, since then at least one Yukawa coupling becomes large (compare the results found in [52]), while for $|\sin 2\,\theta_{R}|$ being maximal the Yukawa couplings and, hence, the size of the cLFV signals are suppressed. This behaviour is very similar for light neutrino masses with IO and $m_{0}=0.015\,\mbox{eV}$ as well as in situations in which the lightest neutrino mass $m_{0}$ is set to zero.
@@ 4.5 Results for Case 3 a)
giving rise to $\sin^{2}\theta_{13}\approx 0.022$ and $\sin^{2}\theta_{23}\approx 0.61$. The solar mixing angle is fixed by the choice of the CP symmetry, i.e. the parameter $s$, and the free angle. All values of $s$, $0\leq s\leq n-1$, allow for a good fit of the lepton mixing angles, usually for two different values of the free angle. For examples of viable fits, see [36] and more recently [52].
@@
For the choice $m$ even and $s$ odd (for $n=34$) we expect the results to depend on $\theta_{R}$ and this dependence should be similar to the one found for Case 1), compare [52]. As can be observed in fig. 13, the dependence is much milder (for light neutrino masses with NO and $m_{0}=0.03\,\mbox{eV}$) than in the other studied cases, Case 1) (see fig. 3), Case 2) for $t$ odd (see fig. 10) and Case 3 b.1) for $m$ even and $s$ odd or vice versa (see fig. 19). We have confirmed that for $m$ odd and $s$ even (employing $n=16$) a very similar result is obtained. Assuming light neutrino masses with IO and $m_{0}=0.015\,\mbox{eV}$ also leads to plots similar to those shown in fig. 13. For NO light neutrino masses and $m_{0}$ being very small ($m_{0}=0$ or $m_{0}=10^{-5}\,\mbox{eV}$) we find instead a larger dependence on $\theta_{R}$ like for Case 1) in fig. 3, when considering e.g. $n=16$, $m=1$ and $s=2$. Regarding the experimental prospects to detect cLFV signals involving the muon and the electron for Case 3 a) and $m$ even and $s$ odd or vice versa, these are similar to those for $m$ and $s$ both even or both odd, namely for the studied examples we obtain $4\times 10^{-21}\lesssim\mathrm{BR}(\mu\to e\gamma)\lesssim 6\times 10^{-16}$, $\mathrm{BR}(\mu\to 3\,e)\gtrsim 10^{-22}$ and $\mathrm{CR}(\mu-e,\mathrm{Al})\gtrsim 10^{-21}$, if the cancellation region, see footnote 20, is avoided.
Lastly, we would like to compare the different choices of $n$, $m$ and $s$ that we have analysed numerically. Like for Case 2), see fig. 11, we make use of averages corresponding to a certain binning of the data in the parameter $\mu_{0}$. We find that the results of the BRs and CR are nearly independent of the choice of $n$, $m$ and $s$, as can be clearly seen in fig. 14, where we show the results for light neutrino masses with NO and non-zero $m_{0}$. This behaviour can be explained with the fact that (almost) no enhancement of the BRs and CR is observed for certain values of the angle $\theta_{R}$, compare fig. 13. The corresponding plots for light neutrino masses with IO and $m_{0}=0.015\,\mbox{eV}$ reveal the same features, as expected from the above discussion.
@@
Figure 15: Case 3 b.1), $m$ and $s$ both even. Predictions for $\mathrm{BR}(\mu\to e\gamma)$, $\mathrm{BR}(\mu\to 3\,e)$ and $\mathrm{CR}(\mu-e,\mathrm{Al})$ in the $\frac{s}{n}-\theta_{L}$-plane in the upper, middle and lower row. The group theory parameters are chosen as $n=20$ and $m=10$. The conventions are like in fig. 5 of Case 2) and $t$ even ($u$ even). Simplified versions of these plots, with only contour lines or only experimentally preferred regions, can be found in appendix E, see figs. 24 and 25. Plots reflecting the features of examples with $m$ and $s$ both odd can be found in figs. 16 and 17, since these are displayed for the choice $\theta_{R}=0$.
@@
Figure 16: Case 3 b.1), $m$ odd, $s$ even and $\theta_{R}=0$. Predictions for $\mathrm{BR}(\mu\to e\gamma)$, $\mathrm{BR}(\mu\to 3\,e)$ and $\mathrm{CR}(\mu-e,\mathrm{Al})$ in the $\frac{s}{n}-\theta_{L}$-plane in the upper, middle and lower row. The group theory parameters are chosen as $n=20$ and $m=9$. For the remaining conventions see fig. 15. Simplified versions of these plots, with only contour lines or only experimentally preferred regions, can be found in appendix E, see figs. 26 and 27. The dependence on $\theta_{R}$ is illustrated with six different values of $\theta_{R}$ in different colours in fig. 30 in the same appendix. Results for an example of the combination $m$ even and $s$ odd and different values of $\theta_{R}$ are found in fig. 32 also in this appendix.
@@
Figure 17: Case 3 b.1), $m$ odd, $s$ even and $\theta_{R}=0$. Predictions for $\mathrm{BR}(\mu\to e\gamma)$, $\mathrm{BR}(\mu\to 3\,e)$ and $\mathrm{CR}(\mu-e,\mathrm{Al})$ in the $\frac{s}{n}-\theta_{L}$-plane in the upper, middle and lower row. The group theory parameters are chosen as $n=20$ and $m=11$. For the remaining conventions see fig. 15. Simplified versions of these plots, with only contour lines or only experimentally preferred regions, can be found in appendix E, see figs. 28 and 29. The dependence on $\theta_{R}$ is illustrated with six different values of $\theta_{R}$ in different colours in fig. 31 in the same appendix. Results for an example of the combination $m$ even and $s$ odd and different values of $\theta_{R}$ are found in fig. 32 also in this appendix.
@@
We begin the analysis of the BRs and CR by evaluating their sizes in the $\frac{s}{n}-\theta_{L}$-plane [^23] The ratio $\frac{s}{n}$ is taken to be continuous, $0\leq\frac{s}{n}\leq 1$, and $s$ is either assumed to be even or odd. for a fixed ratio of $\frac{m}{n}$, $\frac{m}{n}=\frac{9}{20}$, $\frac{m}{n}=\frac{10}{20}$ and $\frac{m}{n}=\frac{11}{20}$, as well as for fixed values of the two scales $\mu_{0}$ and $M_{0}$, $\mu_{0}=1\,\mbox{keV}$ and $M_{0}=3\,\mbox{TeV}$, see figs. 15-17. The colour-coding in these figures is the same as in fig. 5 for Case 2). In particular, the dark (light) grey regions in the plots for $\mathrm{CR}(\mu-e,\mathrm{Al})$ indicate the parameter space in which the lepton mixing angles are accommodated to a certain degree, i.e. $\chi^{2}\leq 100\,(300)$. These regions are found to be consistent with those shown in corresponding plots in [36, 52]. The plots in the left in figs. 15-17 assume that light neutrino masses follow NO with $m_{0}=0.03\,\mbox{eV}$, while the plots in the right in these figures are based on an IO light neutrino mass spectrum with $m_{0}=0.015\,\mbox{eV}$. Fig. 15 represents the combination $m$ and $s$ both even and, thus, no dependence on $\theta_{R}$ is expected, and also observed, while in figs. 16 and 17 $m$ is odd and $s$ even such that the results depend on $\theta_{R}$. Here, they are shown for $\theta_{R}=0$, for which the matrix combination in square brackets in eq. (26) is diagonal. We complete the set of plots by figs. 30-32 in appendix E in which we display for $\frac{m}{n}=\frac{9}{20}$ and $\frac{m}{n}=\frac{11}{20}$ and $s$ even as well as for $\frac{m}{n}=\frac{10}{20}$ and $s$ odd the results for six different values of the angle $\theta_{R}$ in different colours, similar to e.g. fig. 6 for Case 2) and $t$ odd. We note that the results for $\frac{m}{n}=\frac{9}{20}$ ($\frac{m}{n}=\frac{11}{20}$) and $s$ even agree for the choice $\theta_{R}=0$ with those for $\frac{m}{n}=\frac{9}{20}$ ($\frac{m}{n}=\frac{11}{20}$) and $s$ odd and, similarly, the choice $\frac{m}{n}=\frac{10}{20}$ and $s$ odd with $\theta_{R}=0$ reproduces the results of $\frac{m}{n}=\frac{10}{20}$ and $s$ even. Consequently, the results for all possible combinations of $m$ and $s$ even and odd are covered with the six figures, figs. 15-17 and figs. 30-32. We see for $m$ and $s$ both even or both odd as well as for $m$ even and $s$ odd or vice versa and $\theta_{R}=0$ that in most situations the prospective bound on $\mathrm{BR}(\mu\to e\gamma)$ is passed in the entire parameter space – at least at the $3\,\sigma$ level, whereas the limit expected from Mu3E Phase 1 excludes about half of the parameter space at the $3\,\sigma$ level and the future bound from Mu3E Phase 2 considerably reduces the viable regions of parameter space. Employing the expected limits from COMET and Mu2e only leaves a rather small portion of parameter space in the $\frac{s}{n}-\theta_{L}$-plane allowed, which in general is hardly compatible with the regions preferred by the measured values of the lepton mixing angles. In addition, we observe that the future bounds are slightly more likely to be satisfied for light neutrino masses with NO and $m_{0}=0.03\,\mbox{eV}$ than in the case of an IO light neutrino mass spectrum with $m_{0}=0.015\,\mbox{eV}$ (compare plots in the left with those in the right in figs. 15-17).
When considering the dependence on $\theta_{R}$ for the combinations $m$ even and $s$ odd or vice versa, see especially figs. 30-32 in appendix E, we note that the shape and size of the parameter space passing a certain future bound depends on the actual value of $\theta_{R}$ and for the larger ones among the shown values it is usually smaller. Furthermore, the regions of parameter space that lead to an agreement of the lepton mixing angles with experimental data shift as function of $\theta_{R}$. It is, thus, necessary to investigate each value of $\theta_{R}$ separately in order to make more quantitative statements about the constraints imposed by the future limits on the different $\mu-e$ transitions and their compatibility with the restrictions coming from fitting the lepton mixing angles well.
@@
Again, we have fixed the two scales $\mu_{0}$ and $M_{0}$ to $\mu_{0}=1\,\mbox{keV}$ and $M_{0}=3\,\mbox{TeV}$, respectively. From these estimates we see that the future bound on $\mathrm{BR}(\mu\to e\gamma)$ is usually passed as well as the limit on $\mathrm{BR}(\mu\to 3\,e)$ expected from Mu3E Phase 1, whereas Mu3E Phase 2 can exclude this scenario and also the experiments COMET and Mu2e have this potential, since $\mathrm{CR}(\mu-e,\mathrm{Al})$ is larger than $2.7\times 10^{-16}$ for the chosen values of the scales $\mu_{0}$ and $M_{0}$ and the assumed light neutrino mass spectra.